Presentation is loading. Please wait.

Presentation is loading. Please wait.

27 September 2005A.Ciocio - ID Week1 SCT Service Installation Overview Status of Type IV Cable Installation Type II/III Cable Installation Plan for the.

Similar presentations

Presentation on theme: "27 September 2005A.Ciocio - ID Week1 SCT Service Installation Overview Status of Type IV Cable Installation Type II/III Cable Installation Plan for the."— Presentation transcript:

1 27 September 2005A.Ciocio - ID Week1 SCT Service Installation Overview Status of Type IV Cable Installation Type II/III Cable Installation Plan for the next six months Safety course today @ 12:45

2 27 September 2005A.Ciocio - ID Week2 Overview 2044 (+spares) double length Type IV cables (Power & Control) from PP3 racks through holes to US(A)15 racks 4088 (+100 spares) Type II/III (spliced on the outside of the calorimeter and up at z=0) from PPB(F)1 to PP3 racks DCS, Heater Cables, and PP3 (will be discussed in the ID installation) 8 PP3 racks in each side of the cavern (US&USA sides) –Up to 10 crates of 30 PP3 units 11 power racks in US15 and 11 in USA15 –Up to 4 crates of 48 channels TC installation team: –S.Malyukov, E.Kudinov, L.Reed, N.Massol –27 people from IHEP Protvino –Several teams are working in parallel (~9 people per team)

3 27 September 2005A.Ciocio - ID Week3 PP3 Racks in UX15 (US15 side) Y.52-4.X7, Y.53-4.X7 Y.52-1.X6, Y.46-1.X6 Y.43-1.X6, Y.34-1.X6 Y.51-2.X2, Y.52-2.X2

4 27 September 2005A.Ciocio - ID Week4 PP3 Racks in UX15 (USA15 side) Y.51-23.X8 Y.52-23.X8 Y.53-23.X8 Y.51-26.X5, Y.53-26.X5 Y.52-25.X2, Y.53-25.X2 Y.51-25.X1

5 27 September 2005A.Ciocio - ID Week5 Layout of PP3 rack in UX cavern

6 27 September 2005A.Ciocio - ID Week6 PP3 to US PP3 rackUS rackhole UX/US Y.37-1.X6Y.05-14.S212 Y.43-1.X6Y.10-15.S211 Y.43-1.X6Y.11-15.S210 Y.47-1.X6Y.11-15.S210 Y.43-1.X6Y.12-15.S29 Y.47-1.X6Y.12-15.S29 Y.52-1.X6Y.23-15.S27 Y.52-4.X7Y.23-15.S27 Y.52-4.X7Y.24-15.S221 Y.53-4.X7Y.25-15.S220 Y.53-2.X2Y.26-15.S26 Y.51-2.X2Y.27-15.S219 Y.53-4.X7Y.27-15.S219 Y.51-2.X2Y.28-15.S25 Y.52-1.X6Y.28-15.S25 Y.52-2.X2Y.29-15.S24

7 27 September 2005A.Ciocio - ID Week7 Installation of Type IV Cables Type IV cables are delivered in double lengths and pulled between the UX15 and the US(A) caverns ~800 pieces (double length cables) installed on US side (almost completed) –Up to 40 cables per day pulled from UX to US(A) UX side US(A ) side PP3Power Supply Dummy PP3 Racks on Level 6 US side Hole

8 27 September 2005A.Ciocio - ID Week8 In US15 Cut cables and connectors on the power supply side installed in situ First rack (200 cables/connectors) completed early in July –Work performed by SEDELEC (subcontract of KontaktSysteme that provides the double length cables –Many mistakes (~20 connectors have just been repaired) several label mismatch still to be fixed –They raised the cost significantly for future connector work New Plan –TC will help with connector attachment –Nicolas Massol joined TC in September. One of his tasks is to organize connector activity –Work with Petr Kubik (Prague) to discuss technical aspects and qualify samples made by other companies being contacted –Meanwhile cables are being cut to some extra length and ~ all tooling has been purchased to make prototypes

9 27 September 2005A.Ciocio - ID Week9 Delivery Issues Installation rate driven by delivery schedule –Delivery in July/August was very slow Vacation and QA problems Connectors quality Bobbins damage –Bobbins already at CERN are going to be sent back to the fabrication company (Novacavi) to get cleaned up Delay of components delivery from CERN to company Need to keep contact with CERN purchasing and Vendor Ramp up –200 pieces/week expected soon after bobbins damage crisis is solved but so far the max number of pieces we have seen is 120 –Hopefully all type-IV delivered by mid-November

10 27 September 2005A.Ciocio - ID Week10 Delivery (QA - Connectors) Connector’s signal wires too long Some cables couldn’t be used One week halt of production

11 27 September 2005A.Ciocio - ID Week11 Delivery (QA - Packaging) Packaging: palette were sent with only plastic wrap around bundles Cables were getting wet during transportation and delivery at CERN About 40 cables lost (~30 in one delivery)

12 27 September 2005A.Ciocio - ID Week12 Delivery (QA - Bobbins) Cable bobbins dirty and damaged Bobbins are kept outside (at CERN) for long time Other (than water) damage under investigation Company claims delay in fabrication to deal with these problems Don’t know yet delay due to shipping bobbins back to Italy

13 27 September 2005A.Ciocio - ID Week13 Type II/III Cables Double harness ordered from Ineo Work on splicing started in August Installation of connectors on the PP3 end at the factory –Need to provide exact length for both types –Use dummy PP1 with some load to test that the splice didn’t get damaged after installation –JST and FCI connectors very delicate – better to test at the PP3 end Two Contracts first to produce all the harnesses and then to install the connectors Splice is verified at the company Need QA going on during production Last visit by Pepe B. Maarit W. and Laurent Gillain(CERN purchasing) quite satisfactory By November all harness will be made Cables with PP3 connectors between November and March

14 27 September 2005A.Ciocio - ID Week14 Current Installation Activities (Type IV) Overall coordination (Sandra) –Work closely with Sergei Malyukov and the TC team –Provide complete layout/mapping for racks in both sides –Bjarne is still helping from Bergen –Regular presence to check mapping/labels and prepare racks –Discussion on installation priority and occasional revise original layout –update DB table and EDMS documents* –Organize cable testing shifts Connector at the power supply end (Petr Kubik/Prague) Cable testing (Sandra, Petr and SCT Collaboration) Follow up cable procurement and delivery (Pippa, Sandra, Maarit/TC) PP3 installation soon (Ashley French/Melbourne with Petr) * SCT Cables EDMS ID: 624679, Cable Procurement Follow up table ID:508310, Overall Safety and Health Procedure ID: 593278

15 27 September 2005A.Ciocio - ID Week15 Type IV Testing Tester designed and provided by Ljubljana (Gregor, Erik) –2 Test box + 2 Laptops + test cables –Gregor and Erik upgraded features of tester in July –Web based DB Test steps: –Cables are tested at reception (acceptance test) –After being pulled from UX15 to US(A)15 (uncut-double length) –After they are cut in the US side at connectors are installed Shift planner (developed by D. Robinson) works very well for planning the activities and have people to sign up –(register and contact me) What is needed to be able to take a shift: 1)take the 3-levels safety training course (at the pompiers station) 2)make the access request on EDH (ATLAS-U, LHC Tunnel) 3)get your safety boots & helmet at the CERN store (or borrow them)

16 27 September 2005A.Ciocio - ID Week16 Procurement Schedule Type IV delivered and installed December Connector work of Type IV (power supply end) could start in 3 weeks and it will take ~4 months January Type II/III double harness with splicingJanuary single cable with PP3 connector Nov-March PP3 delivery and installationOct/Nov

17 27 September 2005A.Ciocio - ID Week17 Work Plan

18 27 September 2005A.Ciocio - ID Week18 Safety Training Course Building 65-1-003 (Pompiers station) At 12:45 Please arrive 5 minutes earlier to register

Download ppt "27 September 2005A.Ciocio - ID Week1 SCT Service Installation Overview Status of Type IV Cable Installation Type II/III Cable Installation Plan for the."

Similar presentations

Ads by Google